# Bidirectional Hierarchical Protein Multi-Modal Representation Learning

### v1

@@ 3.2.1 Single-protein representation task
ScanNet $\times$ 0.732 0.712 0.735 0.605 0.694
ProNet (Amino Acid) $\times$ 0.817 0.705 0.691 0.577 0.685
Serial fusion $\checkmark$ 0.807 0.728 0.726 0.592 0.700
Local Bi-Hierarchical Fusion $\checkmark$ 0.839 0.755 0.714 0.594 0.716
Global Bi-Hierarchical Fusion $\checkmark$ 0.828 0.733 0.735 0.620 0.714
Table 4: Performance assessment for predicting protein-protein binding sites (PPBs) is presented with the Area Under the Curve for the Precision-Recall (AUCPR) metric. The proteins in the test set are categorized into four distinct, non-overlapping groups. For the masif-site, only aggregated performance is displayed, as its training dataset differs from ours. Entries in bold highlight the top performance. The top two results are highlighted as 1st and 2nd.
3.2.2 Protein-molecules representation task
Ligand Binding affinity. Protein-ligand binding affinity (LBA) prediction is a critical task in drug discovery, as it estimates the interaction strength between a candidate drug molecule and a target protein. For this study, we utilize the PDBbind database [20], a curated resource of protein-ligand complexes sourced from the Protein Data Bank (PDB), annotated with their respective binding strengths. To ensure a robust evaluation, the dataset is partitioned such that proteins in the test set share no more than 30% or 60% sequence identity with any protein in the training set.
The advancement of interaction-free methods as illustrated in  Fig. 3 for binding affinity prediction emphasizes the need for sophisticated computational techniques independent of physical interaction data. These methods employ separate neural networks and representations for proteins and molecules, with proteins providing a more complex computational canvas due to their intricate structures and dynamic functions. In our study, we represent proteins using ProNet and ligands as outlined in [19, 16]. We evaluate these methods at both the amino-acid and all-atom levels for a comprehensive assessment of binding affinity prediction. Our evaluations involve comparing root mean squared error (RMSE), Pearson correlation ($R_{p}$), and Spearman correlation ($R_{s}$), with the main results detailed in Table 1. ProNet, previously the state-of-the-art, serves as our baseline. Serial fusion improves upon this, reducing RMSE by 0.03-0.05. Our Bi-Hierarchical Fusion approaches further decrease RMSE by 0.01-0.08, with global Bi-Hierarchical Fusion achieving the lowest RMSE in the challenging 30% sequence identity split, demonstrating strong generalizability.
3.2.3 Protein-protein representation task
@@
For this task, we evaluate PPBS using the ScanNet metric [17] based on a detailed dataset from the Dockground database, which includes 20,000 protein chains and spans various complex types, covering 5 million amino acids, with 22.7% identified as PPBS. We assess performance using the area under the precision-recall curve (AUCPR). Our test set proteins, aligned with the ScanNet setup, are categorized into four exclusive groups: (a) Test 70%: Proteins sharing at least 70% sequence identity with training set examples. (b) Test homology: Proteins with up to 70% sequence identity to any training example but within the same protein superfamily (H-level in CATH classification). (c) Test topology: Proteins sharing similar topology (T-level in CATH classification) but not the same superfamily. (d) Test none: Proteins that do not fit any previous categories. Our results, detailed in Table 4, show that while ProNet was our initial benchmark and typically underperformed relative to ScanNet, employing a pretrained protein large language model via Bi-Hierarchical Fusion markedly enhanced performance on all baseline metrics. Specifically, local Bi-Hierarchical Fusion exceeds ScanNet in most metrics, and global Bi-Hierarchical Fusion outperforms ScanNet in all metrics.
@@
In this study, we adopted the dataset configuration from Scannet [17], and extracted data from the SabDab database [7]. This dataset includes 3,756 protein chains, each annotated with BCEs, where 8.9% of the residues are identified as BCEs—a figure that likely underestimates the actual percentage. The dataset was segmented into five subsets for cross-validation, with each pair of sequences from different subsets having no more than 70% sequence identity. Table 5 illustrates the performance evaluation of BCE prediction using the Area Under the Precision-Recall Curve (AUCPR). The results indicate that both the local and global Bi-Hierarchical Fusion approaches surpass the serial fusion method in effectiveness. Notably, the Global Bi-Hierarchical Fusion approach achieved a score of 0.2418, significantly outperforming all other baselines, including the current state-of-the-art, ScanNet, by a considerable margin.
@@ Acknowledgements
This work is supported by the RadBio-AI project (DE-AC02-06CH11357), U.S. Department of Energy Office of Science, Office of Biological and Environment Research, the Improve project under contract (75N91019F00134, 75N91019D00024, 89233218CNA000001, DE-AC02-06-CH11357, DE-AC52-07NA27344, DE-AC05-00OR22725), the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration.
@@ A.2 Hyperparameters and architectures
  • Gaussian noise: if True, will add noise to the node features before each interaction block [19].
  • Euler noise: if True, will add noise to Euler angles [19]
A.3 Dataset Description
@@
A.3.1 Reaction classification
Enzymes, which act as biological catalysts, are classified by Enzyme Commission (EC) numbers according to the reactions they catalyze. To evaluate our approach, we adopt the dataset and experimental protocol introduced by Hermosilla et al. [10]. The dataset is divided into 25670 proteins for training, 2852 for validation, and 5598 for testing. Each EC number is present across all three splits, and protein chains sharing more than 50% sequence similarity are grouped together.
A.3.2 Model quality assessment
The Critical Assessment of Structure Prediction (CASP) [4] is a long-running international competition focused on protein structure prediction, with CASP14 being the latest edition. In this challenge, newly resolved experimental structures are withheld to evaluate predictive performance. Following the protocol of [21], we divide the decoy sets by target and release year: CASP11-12 are used for training, CASP13 for validation, and CASP14 for testing. This results in a train/validation/test split of 22885/3536/6492.
A.3.3 Ligand binding affinity
PDBbind includes X-ray crystal structures of proteins complexed with small molecules and peptide ligands. We utilize a dataset derived from PDBbind [20], which is available in two variants based on sequence identity thresholds of 30% and 60%. The 30% identity split yields train/validation/test sets of 2929/325/480, while the 60% identity split produces train/validation/test sets of 3016/335/424.
A.3.4 Protein-protein binding sites
For the tasks of protein–protein binding site (PPBS) prediction and B-cell conformational epitope (BCE) prediction, we utilize the dataset provided by ScanNet [17]. As described in Section 3.2.3, ScanNet defines four distinct test groups for PPBS evaluation. Instead of using separate validation sets for each group as defined in ScanNet, we derive a single validation set from the training partition. This results in 11225 proteins for training, 1247 for validation, and the following counts for testing: 539 for Test 70%, 1453 for Test Homology, 893 for Test Topology, 1050 for Test None, and 3935 for Test All.

### v2

@@ 3.2.3 Protein-protein representation task
ScanNet $\times$ 0.732 $\pm$ 0.005 0.712 $\pm$ 0.004 0.735 $\pm$ 0.005 0.605 $\pm$ 0.006 0.694 $\pm$ 0.003
ProNet (Amino Acid) $\times$ 0.817 $\pm$ 0.007 0.705 $\pm$ 0.004 0.691 $\pm$ 0.006 0.577 $\pm$ 0.003 0.685 $\pm$ 0.002
Serial fusion $\checkmark$ 0.807 $\pm$ 0.003 0.728 $\pm$ 0.002 0.726 $\pm$ 0.004 0.592 $\pm$ 0.002 0.700 $\pm$ 0.003
Local Bi-Hierarchical Fusion $\checkmark$ 0.839 $\pm$ 0.004 0.755 $\pm$ 0.003 0.714 $\pm$ 0.002 0.594 $\pm$ 0.003 0.716 $\pm$ 0.002
Global Bi-Hierarchical Fusion $\checkmark$ 0.828 $\pm$ 0.004 0.733 $\pm$ 0.002 0.735 $\pm$ 0.003 0.620 $\pm$ 0.002 0.714 $\pm$ 0.003
Table 4: Performance assessment for predicting protein-protein binding sites (PPBs) is presented with the Area Under the Curve for the Precision-Recall (AUCPR) metric. The proteins in the test set are categorized into four distinct, non-overlapping groups. For the masif-site, only aggregated performance is displayed, as its training dataset differs from ours. Entries in bold highlight the top performance. The top two results are highlighted as 1st and 2nd. The results for ScanNet are taken directly from the original paper (Tubiana et al., 2022), which reports performance across 10 random seeds. For our methods, we report the mean and standard deviation over 5 independent runs.
@@
For this task, we evaluate PPBS using the ScanNet metric (Tubiana et al., 2022) based on a detailed dataset from the Dockground database, which includes 20,000 protein chains and spans various complex types, covering 5 million amino acids, with 22.7% identified as PPBS. We assess performance using the area under the precision-recall curve (AUCPR). Our test set proteins, aligned with the ScanNet setup, are categorized into four exclusive groups: (a) Test 70%: Proteins sharing at least 70% sequence identity with training set examples. (b) Test homology: Proteins with up to 70% sequence identity to any training example but within the same protein superfamily (H-level in CATH classification). (c) Test topology: Proteins sharing similar topology (T-level in CATH classification) but not the same superfamily. (d) Test none: Proteins that do not fit any previous categories. Our results, detailed in Table 4, show that while ProNet was our initial benchmark and typically underperformed relative to ScanNet, employing a pretrained protein large language model via Bi-Hierarchical Fusion markedly enhanced performance on all baseline metrics. Specifically, local Bi-Hierarchical Fusion exceeds ScanNet in most metrics, and global Bi-Hierarchical Fusion outperforms ScanNet in all metrics.
@@
In this study, we adopted the dataset configuration from Scannet (Tubiana et al., 2022), and extracted data from the SabDab database (Dunbar et al., 2014). This dataset includes 3,756 protein chains, each annotated with BCEs, where 8.9% of the residues are identified as BCEs—a figure that likely underestimates the actual percentage. The dataset was segmented into five subsets for cross-validation, with each pair of sequences from different subsets having no more than 70% sequence identity. Table 5 illustrates the performance evaluation of BCE prediction using the Area Under the Precision-Recall Curve (AUCPR). The results indicate that both the local and global Bi-Hierarchical Fusion approaches surpass the serial fusion method in effectiveness. Notably, the Global Bi-Hierarchical Fusion approach achieved a score of 0.2418, significantly outperforming all other baselines, including the current state-of-the-art, ScanNet, by a considerable margin.
@@ Acknowledgements
This work is supported in part by the RadBio-AI project (DE-AC02-06CH11357), U.S. Department of Energy Office of Science, Office of Biological and Environment Research, the Improve project under contract (75N91019F00134, 75N91019D00024, 89233218CNA000001, DE-AC02-06-CH11357, DE-AC52-07NA27344, DE-AC05-00OR22725), the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration.
@@ A.2 Hyperparameters and architectures
  • Gaussian noise: if True, will add noise to the node features before each interaction block (Wang et al., 2023).
  • Euler noise: if True, will add noise to Euler angles (Wang et al., 2023)
A.3 Ablation Study: Number of ESM layers
ESM Layer # of Parms MSE $\downarrow$ $R_{p}$ $\uparrow$ $R_{s}$ $\uparrow$
6 8M 1.377 0.736 0.741
12 35M 1.365 0.739 0.736
30 150M 1.323 0.770 0.761
Table 7: Binding Affinity 60% with Local Bi-Hierarchical Fusion. Ablation study showing that increasing ESM layers improves performance
A.4 Dataset Description
@@
A.4.1 Reaction classification
Enzymes, which act as biological catalysts, are classified by Enzyme Commission (EC) numbers according to the reactions they catalyze. To evaluate our approach, we adopt the dataset and experimental protocol introduced by Hermosilla et al. (2020). The dataset is divided into 25670 proteins for training, 2852 for validation, and 5598 for testing. Each EC number is present across all three splits, and protein chains sharing more than 50% sequence similarity are grouped together.
A.4.2 Model quality assessment
The Critical Assessment of Structure Prediction (CASP) (Cheng et al., 2019) is a long-running international competition focused on protein structure prediction, with CASP14 being the latest edition. In this challenge, newly resolved experimental structures are withheld to evaluate predictive performance. Following the protocol of (Wu et al., 2023), we divide the decoy sets by target and release year: CASP11-12 are used for training, CASP13 for validation, and CASP14 for testing. This results in a train/validation/test split of 22885/3536/6492.
A.4.3 Ligand binding affinity
PDBbind includes X-ray crystal structures of proteins complexed with small molecules and peptide ligands. We utilize a dataset derived from PDBbind (Wang et al., 2005), which is available in two variants based on sequence identity thresholds of 30% and 60%. The 30% identity split yields train/validation/test sets of 2929/325/480, while the 60% identity split produces train/validation/test sets of 3016/335/424.
A.4.4 Protein-protein binding sites
For the tasks of protein–protein binding site (PPBS) prediction and B-cell conformational epitope (BCE) prediction, we utilize the dataset provided by ScanNet (Tubiana et al., 2022). As described in Section 3.2.3, ScanNet defines four distinct test groups for PPBS evaluation. Instead of using separate validation sets for each group as defined in ScanNet, we derive a single validation set from the training partition. This results in 11225 proteins for training, 1247 for validation, and the following counts for testing: 539 for Test 70%, 1453 for Test Homology, 893 for Test Topology, 1050 for Test None, and 3935 for Test All.